### v1

# Disentangling quantum autoencoder

Abstract
Entangled quantum states are highly sensitive to noise, which makes it difficult to transfer them over noisy quantum channels or to store them in quantum memory. Here, we propose the disentangling quantum autoencoder (DQAE) to encode entangled states into single-qubit product states. The DQAE provides an exponential improvement in the number of copies needed to transport entangled states across qubit-loss or leakage channels compared to unencoded states. The DQAE can be trained in an unsupervised manner from entangled quantum data. We demonstrate variational quantum algorithms for general states, and a Metropolis algorithm for stabilizer states. For particular classes of states, the number of training data needed to generalize is surprisingly low: For stabilizer states, DQAE generalizes by learning from a number of training data that scales linearly with the number of qubits, while only $1$ training data is sufficient for states evolved with the transverse-field Ising Hamiltonian. Our work provides practical applications for enhancing near-term quantum computers.
I Introduction
Entanglement is a key feature of quantum mechanics that describes the non-local quantum correlations, and is a necessary ingredient for achieving quantum advantages in quantum communication, quantum sensing and quantum computing [1]. However, at the same time, entanglement makes quantum states fragile, as the non-local correlations are highly susceptible to noise that affects quantum states in experiment. In particular, even noise that acts only on a local subsystem of the entangled state can already collapse the global wavefunction into a classical mixture of states without useful quantum correlations.
To make states robust against noise, various methods have been proposed, such as quantum error correction where the logical information is encoded redundantly into many physical qubits [2]. However, quantum error correction is highly resource-intensive, and not possible in noisy intermediate-scale quantum computers [3]. Furthermore, error correction may be infeasible in scenarios where physical constraints exist, such as a limited number of qubits or long-distance communication between parties.
Alternatively, can reducing the entanglement of states enhance their robustness to noise?? In classical settings, reducing non-local correlations to simplify data has proven to be a fruitful approach. For example, in classical neural networks, variational autoencoders have been used to disentangle highly correlated information into factorized probability distributions [4, 5]. This has allowed to find simpler representations of complicated data that can enhance machine learning.
In quantum information, quantum autoencoders have been proposed to compress high-dimensional quantum states into a lower-dimensional subspace [6, 7, 8]. They can finding a more compact representation of quantum states [9, 10, 11, 12] to enhance information processing. Notably, quantum autoencoders can be used to denoise quantum states [13, 14, 15], and discover new quantum error correction codes [16]. They have been applied to enhance quantum communication in photonic systems [17]. Quantum autoencoders have been fruitful in experiment, with several experimental demonstrations [18, 15, 17].
Quantum autoencoders have been implemented using variational quantum algorithms [19], in which the unitary is represented by a parameterized quantum circuit that is trained on a few example states. To train the circuit parameters, the encoding unitary is learned [20, 21]. The training can be challenging, requiring one to avoid exponential small gradients in barren plateaus [22]. Further, to successfully converge to a good solution, one often requires Overparameterization [23, 24, 25, 26, 27, 28]. Additionally, one would like to generalize, i.e. that from a few training states drawn from some distribution, the trained circuit is able to perform well on arbitrary states drawn from the distribution [29, 30, 31, 32, 33, 28].
However, the literature on quantum autoencoders mainly focuses on reducing the number of qubits, i.e. finding a lower-dimensional representation of quantum states. Compressing information to less qubits implies that quantum information is encoded into non-local correlations, strongly increasing the amount of entanglement. Thus, the current types of quantum autoencoder produces highly entangled states, which are sensitive to noise, and have a complicated representation.
Here, we propose the disentangling quantum autoencoder (DQAE). It transforms entangled $N$-qubit quantum states into a tensor products of $N$ single-qubit quantum states, removing entanglement and factorizing the quantum correlations. We apply the DQAE for the task of transporting entangled state through a qubit-loss channel. In comparison to unencoded transport, the DQAE has an exponential advantage in the number of copies needed to transfer at least one error-free state. The DQAE can be trained in an unsupervised manner using entangled quantum data via variational quantum algorithms. We show that the gradient can be computed efficiently from simple Pauli measurements or Bell measurements. We show that for structured circuit types, barren plateaus can be absent, allowing for efficient training. We characterize the generalization capability, i.e. whether training on a few states allows the DQAE to disentangle unseen test states, using the data quantum Fisher information metric (DQFIM). For Ising-type evolution, we astonishingly require only $L=1$ training states and $M\propto N$ circuit parameters to generalize. For stabilizer states, we propose a Metropolis algorithm to train a Clifford ansatz that generalizes using only $L\propto N/4$ training states. With our results, quantum information processing with near-term quantum computers can be greatly enhanced.
II Model
The DQAE $U$, as shown in Fig. 1, disentangles a set $W$ of $N$-qubit quantum state $|\psi\rangle\in W$ into a tensor product of $N$ single-qubit states
$U|\psi\rangle=\bigotimes_{j=1}^{N}|\phi_{j}\rangle\,,$ (1)
where $|\phi_{j}\rangle$ are single-qubit states that depend on $|\psi\rangle$. As $U$ is unitary, one can reconstruct the entangled original state via the inverse $U^{\dagger}$
$U^{\dagger}\bigotimes_{j=1}^{N}|\phi_{j}\rangle=|\psi\rangle\,.$ (2)
From simple counting arguments, one can see that a single DQAE $U$ cannot compress the set of all possible $N$-qubit states $W=\mathcal{S}(2^{N})$ into single-qubit states at the same time. In particular, to represent any $N$-qubit state $|\psi\rangle$ as single-qubit tensor products, one requires $O(2^{N})$ single-qubit states [34, 35, 36]. However, efficient compression is possible when we consider the DQAE for compressing a restricted set $W$ of states. We note that while we only consider compression to $N$ qubits, one can also reduce the number of single-qubit product states to $K\leq N$. Here, one can further enhance the capacity of the DQAE by including classical information into the $N-K$ qubits that were reduced as shown in Appendix A.
Refer to captionRefer to caption
Figure 1: Disentangling quantum autoencoder (DQAE) disentangles $N$-qubit state $|\psi\rangle$ into a tensor product of $N$ single-qubit states $|\phi_{j}\rangle$ via unitary $U(\boldsymbol{\theta})$, and can recover the original state again via the inverse $U^{\dagger}(\boldsymbol{\theta})$. The DQAE can be trained in an unsupervised manner from entangled data. For transferring quantum states via a qubit-loss channel, DQAE gives an exponential improvement in number of copies needed.
Refer to captionRefer to caption
Figure 2: Disentangling quantum autoencoder (DQAE) to transfer quantum states across qubit loss channels with exponential advantage. Entangled input $N$-qubit state $|\psi\rangle$ is transformed by $U(\boldsymbol{\theta})$ into $\bigotimes_{j=1}^{N}|\phi_{j}\rangle=U|\psi\rangle$,a tensor product of $N$ single-qubit states $|\phi_{j}\rangle$. The transformed state is then sent through a qubit loss channel $\mathcal{E}$, and stored in a quantum register by the receiver. We repeat this process $R$ copies of $|\psi\rangle$, and finally store $NR$ qubits in the quantum register. Some qubits are lost due to the channel $\mathcal{E}$ (red crosses), which we detected using the leakage detection protocol of Fig. 3. At least one of each single-qubit state survives with high probability when $R\propto\log(N)$. In this case, we collect $k=1,\dots,N$ single-qubit states $|\phi_{j}\rangle$ unaffected by loss (green circle). With entangler $U^{\dagger}$ we then reconstruct the original state $|\psi\rangle=U^{\dagger}(\boldsymbol{\theta})\bigotimes_{j=1}^{N}|\phi_{j}\rangle$.
III DQAE for qubit-loss channels
A key application of the DQAE is to deal with qubit loss, which poses a major challenge in quantum computing and communication. Here, we will see that the DQAE can enable an exponential improvement.
[…]
$\mathcal{E}_{k}(\rho)=(1-q)\rho+q\text{tr}_{\bar{k}}(\rho)\otimes|\bot\rangle\langle\bot|\,,$ (3)
where with probability $q$ is lost. Here, $\text{tr}_{\bar{k}}(\rho)$ traces out the $k$th qubit. We assume loss occurs on all qubits equally with the total channel
$\mathcal{E}(\rho)=\mathcal{E}_{1}\circ\mathcal{E}_{2}\circ\dots\circ\mathcal{E}_{N}(\rho)\,.$ (4)
For a fully entangled state, loss on any qubit collapses the state into a mixture, and it cannot be used for further processing. As simple example, let us consider the $N$-qubit GHZ state
$|\text{GHZ}\rangle=\frac{1}{\sqrt{2}}(|0\dots 0\rangle+|1\dots 1\rangle)\,.$ (5)
Loss of any qubit collapses the state into a $N-1$ qubit mixed state
$|\text{GHZ}\rangle\rightarrow\frac{1}{2}(|0\dots 0\rangle\langle 0\dots 0|+|1\dots 1\rangle\langle 1\dots 1|)\,.$ (6)
In contrast, for a single-qubit product state
$\bigotimes_{j=1}^{N}|\phi_{j}\rangle\rightarrow\bigotimes_{j\neq k}^{N}|\phi_{j}\rangle\,,$ (7)
the channel only traces out the $k$th qubit state $|\phi_{k}\rangle$, while the other $N-1$ qubits remain in a pure state. Thus, product states are more robust, which we will showcase in the following.
III.1 Unencoded transfer
First, let us consider a direct unencoded transfer of the $R$ copies. We apply channel $\mathcal{E}(|\psi\rangle)$ on the pure $N$-qubit entangled state $|\psi\rangle$ and with a chance of $(1-q)^{N}$ no loss occurs on any qubit. The chance of successfully transferring the state has an exponentially small probability. Now we repeat the protocol with $R$ copies of $|\psi\rangle$. The probability of at least one copy successfully being transferred has a failure probability of
$Q_{\text{e}}=(1-(1-q)^{N})^{R}\,.$ (8)
For $N\gg 1$, we find
$R\approx\frac{-\log(Q_{\text{e}})}{(1-q)^{N}}\,,$ (9)
which implies that we require exponentially large $R$ with $N$ for fixed error probability $Q_{\text{e}}$.
III.2 DQAE
Now, we compress the incoming $N$-qubit state $|\psi\rangle$ into a single-qubit product state using the DQAE. Then, we have the $N$ qubit product state $U|\psi\rangle=\otimes_{j=1}^{N}|\phi_{j}\rangle$, where $|\phi_{j}\rangle\in\mathcal{H}^{2}$ is a single qubit state. Product states are robust against loss, as in the event of loss the remaining qubits remain pure. Our protocol for product states is as follows: We have $R$ copies of $U|\psi\rangle$. After applying $\mathcal{E}(U|\psi\rangle)$, some qubits are lost, however we assume that of the single qubit states $|\phi_{j}\rangle$, at least one of each has survived. One can check whether a qubit has survived with the leakage detection protocol [37] as shown in Fig.3, which also has been realized in experiment [38].
Refer to captionRefer to caption
[…]
We now compute the failure probability $Q_{\text{p}}$ of the DQAE. For $N=1$, the probability that at least one qubit survives the channel is $1-q^{R}$. For $N$ qubits, the product protocol has a failure probability
$Q_{\text{p}}=1-(1-q^{R})^{N}\,.$ (10)
Now, we choose $Q_{\text{p}}\ll 1$ and $q^{R}\ll 1$. In this limit, we find the relation
${}R\approx\frac{\log(\frac{Q_{\text{p}}}{N})}{\log(q)}$ (11)
where $R$ increases only logarithmically with $N$ for fixed $Q_{\text{p}}$.
In Fig. 4, we observe an exponential advantage in copies $R$ for DQAE compared to transferring unencoded entangled states. Further, we find that our formulas Eq. (11) and Eq. (9) for $R$ are good approximations of the exact equations.
Refer to captionRefer to caption
Figure 4: Number of copies $R$ of an $N$-qubit needed to successfully transfer at least one copy through qubit loss channel of Eq. (4), where each qubit is lost with probability $q$. We show transfer with DQAE compared with unencoded transfer of the states for different qubit loss rates $q$. Dashed lines are the approximate formulas Eq. (9) for unencoded and Eq. (11) DQAE. We demand that the protocols may only fail with a total failure probability $Q$ of at most $Q\leq 0.01$.
III.3 Fidelity
We now show how to lower bound the fidelity of the protocol of Fig. 2. We have entangled state $|\psi\rangle$ and then run our protocol of Fig. 2. Let us assume that during the protocol we send $N$ copies of disentangled $U|\psi\rangle$ through the loss channel. We assume that for the $k$th copy, we loose all qubits except the $k$th qubit, remaining $\rho_{k}=\text{tr}_{\bar{k}}(|\psi\rangle)$. We can detect which qubits survived using the leakage detection of Fig. 3, and use the reconstruction step to assemble the $N$ remaining qubits $\rho_{k}$ into a single state $\rho=\bigotimes_{k}\rho_{k}$. If the disentangler is optimal, we can perfectly reconstruct the original stat via $|\psi\rangle=U^{\dagger}\rho U$, as each $\rho_{k}$ is pure.
However, what if the disentangler is not working perfectly? An optimal disentangler has $E=\frac{1}{N}\sum_{k=1}^{N}E_{k}(U|\psi\rangle)=1$, i.e. the disentangled state is a tensor product of single-qubit states.
If the disentangler does not optimally disentangle states, it creates an entangled state with $U|\psi\rangle$ with $E<1$. In this case, for the reconstruction step of Fig. 2, tracing out for each individual qubit $k$ results in a mixed state $\rho_{k}=\text{tr}_{\bar{k}}(U|\psi\rangle)$. The reconstructed state is then given by $\rho_{\text{f}}=U^{\dagger}\bigotimes_{k=1}^{N}\text{tr}_{\bar{k}}(U|\psi\rangle)U$. We can compute the fidelity between the original state $|\psi\rangle$ and the reconstructed state as
$F=\langle\psi|\rho_{\text{f}}|\psi\rangle=\langle\psi|U^{\dagger}\bigotimes_{k=1}^{N}(\text{tr}_{\bar{k}}(U|\psi\rangle\langle\psi|U^{\dagger}))U|\psi\rangle\,,$ (12)
which is $F=1$ only when DQAE encoded state $U|\psi\rangle$ is a product state, and is bounded by $2^{-N}\leq F\leq 1$.
IV Training
Next, we discuss how to train the DQAE such that it can disentangle a given distribution of quantum states $W$. The training will be performed in an unsupervised manner, i.e. the quantum data has no labels. We sample $L$ training states $|\psi^{(\ell)}\rangle\in W$ as the training set $S=\{|\psi^{(\ell)}\rangle\}_{\ell=1}^{L}$ from $W$. The training states are then used to train the DQAE. The DQAE is trained by a parameterized unitary $U(\boldsymbol{\theta})$ via $U(\boldsymbol{\theta})|\psi^{(\ell)}\rangle$, where $\boldsymbol{\theta}=(\theta_{1},\dots,\theta_{M})$ is a $M$-dimensional parameter vector. We minimize a cost function $C_{\text{train}}(\boldsymbol{\theta})$ by varying the parameter $\boldsymbol{\theta}$. The cost function is given by the average purity of the reduced single-qubit density matrix
$C_{\text{train}}(\boldsymbol{\theta})$ $=1-\frac{1}{LN}\sum_{\ell=1}^{L}\sum_{k=1}^{N}E_{k}(U(\boldsymbol{\theta})|\psi^{(\ell)}\rangle)\,$ (13)
$E_{k}(|\psi\rangle$ $=\text{tr}(\text{tr}_{\bar{k}}(|\psi\rangle\langle\psi|)^{2})\,,$ (14)
where $\text{tr}_{\bar{k}}(.)$ is the partial trace over all qubits except the $k$th qubit. We have $C_{\text{train}}(\boldsymbol{\theta})=0$ only when $U(\boldsymbol{\theta})|\psi^{(\ell)}\rangle$ is a pure single-qubit product state for all $\ell$, else $C_{\text{train}}(\boldsymbol{\theta})>0$.
We train the DQAE on the training set of $L$ states and gain the trained unitary $U(\boldsymbol{\theta}^{*})$ with parameter $\boldsymbol{\theta}^{*}$. After training, we evaluate the performance of the DQAE on unseen test data sampled from the same distribution $W$ via the test error
$C_{\text{test}}(\boldsymbol{\theta})=1-\underset{|\psi\rangle\sim W}{\mathbb{E}}\left[\frac{1}{N}\sum_{k=1}^{N}E_{k}(U(\boldsymbol{\theta})|\psi\rangle)\right]\,.$ (15)
We say that the DQAE generalizes when it achieves low test error, i.e. $C_{\text{test}}\approx 0$.
We can compute the single-qubit purity needed for the cost function efficiently using two different methods. First, the purity of the reduced density matrix of the $k$th qubit is given by
$E_{k}(|\psi\rangle)=\text{tr}(\text{tr}_{\bar{k}}(|\psi\rangle)^{2})=\frac{1}{2}(1+\sum_{\alpha\in\{x,y,z\}}\langle\psi|\sigma_{k}^{\alpha}|\psi\rangle^{2})$ (16)
where $\sigma_{k}^{\alpha}$ with $\alpha\in\{x,y,z\}$ are the Pauli operators acting on the $k$th qubit. The Pauli expectation can be easily estimated on a quantum computers by measuring the respective eigenbasis of the Pauli operators, which requires measuring in $3$ different basis ($x$,$y$, and $z$).
As alternative approach, we can use Bell measurements on two copies of the state $|\psi\rangle$. We define the Bell state $|\Psi^{-}\rangle=\frac{1}{\sqrt{2}}(|01\rangle-|10\rangle)$, the $2$-qubit identity matrix $I_{2}$ and the projector $\Pi_{k}=I_{2}^{\otimes k-1}|\Psi^{-}\rangle\langle\Psi^{-}|I_{2}^{\otimes N-k}$ onto the Bell state of the $k$th qubit pair. It can be shown that [39, 40, 41]
$E_{k}(|\psi\rangle)=\text{tr}(\text{tr}_{\bar{k}}(|\psi\rangle)^{2})=1-2\langle\psi|\langle\psi|\Pi_{k}|\psi\rangle|\psi\rangle\,.$ (17)
This can be achieved by measuring only in the Bell basis, as the projectors for all $k$ commute.
We now consider a general parameterized quantum circuits of $d$ layers of the form $|\psi(\boldsymbol{\theta})\rangle=\prod_{j=1}^{d}V_{j}(\boldsymbol{\theta}_{n})T_{j}|0\rangle$ with fixed entangling unitary $T_{j}$, parameter vector $\boldsymbol{\theta}$ and parameterized rotations $V_{j}(\theta_{j})=e^{-i\boldsymbol{\theta}_{j}/2\sigma_{j}}$ generated by some Pauli string $\sigma_{j}$. For such circuits, gradients $\partial_{j}E_{k}(|\psi(\boldsymbol{\theta})\rangle)$ can be measured efficiently with the shift-rule [42]. The shift-rule can be directly adapted to purity measurements [41]
$\partial_{j}E_{k}(|\psi(\boldsymbol{\theta})\rangle)=$
$-2(\langle\psi(\boldsymbol{\theta}+\frac{\pi}{2}e_{j})|\langle\psi(\boldsymbol{\theta})|\Pi_{k}|\psi(\boldsymbol{\theta}+\frac{\pi}{2}e_{j})\rangle|\psi(\boldsymbol{\theta})\rangle$
$\langle\psi(\boldsymbol{\theta}-\frac{\pi}{2}e_{j})|\langle\psi(\boldsymbol{\theta})|\Pi_{k}|\psi(\boldsymbol{\theta}-\frac{\pi}{2}e_{j})\rangle|\psi(\boldsymbol{\theta})\rangle)\,,$
where $e_{j}$ is the unit vector for the $j$th entry of the parameter vector $\boldsymbol{\theta}$.
V Overparameterization and generalization
To successfully train the DQAE, we require two conditions: First, training should converge close to a global minimum, i.e. we have a low training error $C_{\text{train}}$. Additionally, the DQAE needs to generalize from $L$ training data to arbitrary test states, i.e. we have a low test error $C_{\text{test}}$. Training cost is known to converge well when the ansatz circuit is overparameterized, i.e. has enough circuit parameters $M$ to explore the full ansatz space [27, 25, 43]. This can be evaluated via the data quantum Fisher information metric (DQFIM) [28]. In particular, the rank of the quantum Fisher information does not increase further when increasing $M$.
Generalization can be similarly evaluated using the DQFIM. An overparameterized model generalizes (assuming no noise) when the rank of the DQFIM does not increase further upon increasing the number $L$ of training states. Note that for $L=1$, the DQFIM is identical to the well known quantum Fisher information metric [44].
Let use first introduce the data state for training set $S_{L}=\{|\psi^{(\ell)}\rangle\}_{\ell=1}^{L}$
$\rho_{L}=\frac{1}{L}\sum_{\ell=1}^{L}|\psi^{(\ell)}\rangle\langle\psi^{(\ell)}|\,.$ (18)
The DQFIM is defined as [28]
$\mathcal{Q}_{nm}(\rho_{L},U)=$ (19)
$4\mathrm{Re}(\mathrm{tr}(\partial_{n}U\rho_{L}\partial_{m}U^{\dagger})-\mathrm{tr}(\partial_{n}U\rho_{L}U^{\dagger})\mathrm{tr}(U\rho_{L}\partial_{m}U^{\dagger}))\,.$
For a given $L$ training states, the maximal rank of the DQFIM is given as
$R_{L}\equiv\max_{M\geq M_{\text{c}}(L),\boldsymbol{\theta}}\text{rank}(\mathcal{Q}(\rho_{L},U(\boldsymbol{\theta})))\,.$ (20)
where $M_{\text{c}}(L)$ is the critical number of circuit parameters where the model overparameterizes. For $M\geq M_{\text{c}}(L)$, training is likely to converge to the global minima. Note that $M_{\text{c}}(L)$ depends on the number $L$ of training data, and in general the circuit parameters needed to overparameterize increase with $L$.
For generalization (i.e. low test error), one has to supply a critical number of training data $L\geq L_{\text{c}}$ during training
$R_{\infty}=R_{L_{\text{c}}}\equiv\max_{L\geq L_{\text{c}}}R_{L}(\rho_{L},U)\,,$ (21)
such that $R_{L}$ does not increase further upon increase of $L$ (for $M\geq M_{\text{c}}(L)$). The maximal possible $R_{L}$ is the maximal rank of the DQFIM $R_{L_{\text{c}}}\equiv R_{\infty}$. The critical number of training states needed to generalize can be approximated as [28]
$L_{\text{c}}\approx 2R_{\infty}/R_{1}\,.$ (22)
Thus, the rank of the DQFIM allows us to determine how many circuit parameters $M_{\text{c}}$ and training states $L_{\text{c}}$ are needed to generalize, which we will later on apply for the DQAE.
Refer to captionRefer to caption
Figure 5: Parameterized unitaries for disentangler unitary $U(\boldsymbol{\theta})$, which are composed of many layers of the following building blocks: a) Generic hardware-efficient ansatz $U_{\text{HE}}$ of parameterized $y$ and $z$ rotations and fixed CNOT gates arranged in nearest-neighbor fashion. b) Fermionic ansatz $U_{\text{fermion}}$ c) Transverse-field Ising inspired ansatz $U_{\text{Ising}}$.
VI Ansatz circuits
We study different types of ansatz circuits as shown in Fig. 5. In Fig. 5a, we have a generic hardware efficient ansatz, constructed from CNOT gates in a 1D nearest-neighbor fashion $W$ combined with layer of parameterized rotations generated from Pauli operators $\sigma^{x}_{j}$, $\sigma^{z}_{j}$ acting on qubit $j$, which is written as
$U_{\text{HE}}(\boldsymbol{\theta})=\prod_{j=1}^{d}[\bigotimes_{k=1}^{N}e^{-i\sigma^{x}_{k}\theta_{2jN+k}}e^{-i\sigma^{z}_{k}\theta_{2jN+N+k}}]W\,.$ (23)
We also consider a free fermionic ansatz [45], which is composed of Givens rotations as shown in Fig. 5b. Additionally, we consider in Fig. 5c a circuit that represents the evolution of the Ising model $H=H_{xx}+hH_{z}$, with $H_{xx}=\sum_{i}\sigma^{x}_{i}\sigma^{x}_{j+1}$ and $H_{z}=\sum_{i}\sigma^{z}_{j}$, where we have nearest-neighbor Pauli $x$ interactions, and local $z$ interactions. The evolution under this model of $d$ layers is generated via
$U_{\text{Ising}}(\boldsymbol{\theta})=\prod_{j=1}^{d}[\exp(-iH_{xx}\theta_{j}^{x})\exp(-iH_{z}\theta_{j}^{z})]\,.$ (24)
Input : Training data $\{|\psi^{(\ell)}\rangle\}_{\ell=1}^{L}$
  Monotonously decreasing temperature function $T(b)$
Output : DQAE $U$
1 $U=I$
[…]
3 while $C>0$ do
4
5      $b=b+1$
6      choose randomly $a\in\{0,1\}$
7      if a=0 then
8             random $i\in\{1,\dots,N-1\}$, $j\in\{i+1,\dots,N\}$
9            $U^{\prime}=\text{CNOT}_{i,j}U$
10       else
11             random $i\in\{1,\dots,N\}$
12            $U_{\text{cliff}}\in\mathcal{C}_{1}$
13            $U^{\prime}=I_{2}^{\otimes i-1}U_{\text{cliff}}I_{2}^{\otimes N-i}U$
14       end if
15
16      $C^{\prime}=1-\frac{1}{LN}\sum_{\ell=1}^{L}\sum_{k=1}^{N}\text{tr}(\text{tr}_{\bar{k}}(U^{\prime}|\psi^{\ell}\rangle)^{2})$
17      uniform random $r\in[0,1)$
18      $p=\text{min}(1,\exp(-(C^{\prime}-C)/T(b)))$
19      if $r\leq p$ then
20             $U=U^{\prime}$
21            $C=C^{\prime}$
22       end if
23
24 end while
Algorithm 1 Learn Clifford DQAE
VII DQAE with stabilizer states
Stabilizer states and Clifford circuits are one of the fundamental building blocks of quantum information and computing [46]. Due to their simpler structure, they have an efficient algorithm to compute many properties, and even their dynamics can be efficiently learned from quantum states [47]. They are constructed from the Clifford gate set: A combination of CNOT, Hadamard and S-gates, which can be efficiently simulated classically [46].
Here, we propose Algorithm 1 to learn to disentangle stabilizer states, which is inspired from Ref. [48] and the Metropolis algorithm. The goal is to build a disentangling unitary, starting from the identity operator. The idea of the algorithm is to randomly apply either a random single-qubit Clifford unitaries $\mathcal{C}_{1}$ or CNOT gate on randomly chosen qubits. This gate is accepted with probability $p=\text{min}(1,\exp(-(C_{\text{p}}-C)/T(b)))$, where $T(b)$ is a temperature function which monotonously decreases with each step $b$, $C_{\text{p}}$ is the cost function with the added gate, and $C$ without. These steps are repeated until the cost function reaches zero. We empirically find that this algorithm nearly always converges, and we find only a weak dependence on performance depending on the chosen $T(b)$.
For stabilizer input data, the quantum part of the algorithm can be efficiently simulated on classical computers [49]. Note that a classical representation of stabilizer states can be efficiently learned on quantum computers [50]. Thus, we can run this algorithm by first learning a classical representation of the training set, then training the disentangling unitary on a classical computer, where the quantum part is simulated classically.
VIII Results
We now study the DQAE numerically for different ansatze.
VIII.1 Data generation
To generate training and test data, we first draw a fixed unitary $V$ from the respective studied ansatz as shown in Fig. 5. Then, for the training data we prepare $L$ random single-qubit product states $|\psi^{(\ell)}\rangle=\bigotimes_{j}|\phi_{j}^{(\ell)}\rangle$, which gives us the training set
$S_{L}=\{V\bigotimes_{j=1}^{N}|\phi_{j}^{(\ell)}\rangle\}_{\ell=1}^{L}\,,$ (25)
which is used to train the circuit. The test set is drawn similarly, which is not used during training and only for testing the trained ansatz. Our choice of $S_{L}$ ensures that there exists an optimal solution with low test and training error. However, this solution may be difficult to find during training when the number of training data $L$ is low, or the circuit does not have enough circuit parameters $M$.
Refer to captionRefer to caption
Figure 6: We show the variance of the gradient for training the DQAE against qubit number $N$. The gradient is computed at random parameters $\boldsymbol{\theta}$ averaged over 10 random instances. We show a generic ansatz $U_{\text{HE}}$, fermionic ansatz $U_{\text{fermion}}$ and Ising model ansatz $U_{\text{ising}}$. Dashed line is fit with $\text{var}(\nabla C_{\text{train}})=cN^{-\gamma}$, where we find $\gamma\approx 3.8$ for $U_{\text{fermion}}$ and $U_{\text{ising}}$.
VIII.2 Trainability
Now, we study the trainability of the DQAE, i.e. whether the gradients for training are sufficiently large [25]. We show the variance of the gradient $\text{var}(\nabla C_{\text{train}})$ against $N$ in Fig. 6 for different ansatz types as detailed in Fig. 5. We find that for generic hardware efficient ansatz, the variance decays exponentially which indicates a barren plateau [22]. For $U_{\text{ising}}$ and $U_{\text{fermion}}$ we find the gradient decays only polynomially, indicating the absence of barren plateaus for these ansatz unitaries, and the circuit can be trained even when the number of qubits is scaled up.
Refer to captionRefer to caption
Figure 7: Training error $C_{\text{train}}$ (solid line) and test error $C_{\text{test}}$ (dashed) for circuit ansatz $U_{\text{HE}}$ as shown in Fig. 5a. We show error as function of number of training epochs for different number of training states $L$ for $M=400$ circuit parameter in the overparameterized regime. We have critical number of training states $L_{\text{c}}=16$ and $N=4$.
Refer to captionaRefer to caption Refer to captionbRefer to caption Refer to captioncRefer to caption
Figure 8: DQAE for disentangling a generic circuit ansatz $U_{\text{HE}}$ as shown in Fig. 5a. We show the final a) training loss $C_{\text{train}}$, b) test loss $C_{\text{test}}$ and c) number of iterations $E$ needed for training to converge. We show $N=4$ qubits, for different number of training states $L$ and circuit parameters $M$. Dashed lines are thresholds for overparameterization via Eq. (20) and generalization via Eq. (22)
VIII.3 Generic ansatz
Next, we study the generic ansatz, shown in Fig. 5a, in more detail. First, we show a typical training trajectory for the YZ-ansatz in Fig. 7 for different number of training states $L$. We find three regimes in the overparameterized regime $M\gg M_{\text{c}}$: For $L\ll L_{\text{c}}$, $C_{\text{train}}$ converges fast, but the test error remains high. For $L\gg L_{\text{c}}$, dataset is overcomplete and both $C_{\text{train}}$ and $C_{\text{test}}$ converge fast. For $L=15\approx L_{\text{c}}$, we observe that $C_{\text{train}}$ takes a long time to converge, with $C_{\text{test}}$ becoming smaller as well over many epochs.
Next, we study the training and test cost for the generic ansatz for different number of circuit parameters $M$ and training data size $L$. find that the training cost in Fig. 8a becomes small when the circuit is about to overparameterized, i.e. $M>M_{\text{c}}(L)$, where the threshold from the DQFIM is indicated as the horizontal-aligned dashed line. Furthermore, as shown in Fig. 8b, the test error decreases significantly once the model generalizes, which occurs when the number of training states exceeds $L_{\text{c}}\approx 2R_{\infty}/R_{1}$, as indicated by the vertical dashed line. We find that around the overparameterization and generalization threshold, the number of epochs needed to train substantially increases as shown in Fig. 8c. In general, for a generic circuit ansatz, both $L_{\text{c}}$ and $M_{\text{c}}(L)$ scale exponentially with the number of qubits $N$. This has been analytically calculated in Ref. [28], where one finds $L_{\text{c}}=2^{N}$ and $M_{\text{c}}(L)=\text{min}(2^{N+1}L-L^{2}-1,4^{N}-1)$.
VIII.4 Ising ansatz
Next, we study models with polynomial Lie algebra, which can generalize with only polynomial resource cost [28]. In particular, in Fig. 9 we show the training and test loss for DQAE applied to evolution with the Ising model $U_{\text{ising}}$ as shown in in Fig. 5c. Remarkable, for $N\geq 6$, we require only
$L_{\text{c}}=1$ (26)
training states to disentangle input states. The critical number of circuit parameters $M\geq M_{\text{c}}$ needed for overparameterization (and thus converging to a global minima) is low, scaling as $M_{\text{c}}=3N-1$, which we compute via the rank of the DQFIM of Eq. (20). Thus, for the Ising model, DQAE works remarkably well, requiring only $O(N)$ circuit parameters and $O(1)$ training data.
Refer to captionaRefer to caption Refer to captionbRefer to caption
Figure 9: DQAE for disentangling Ising model evolution $U_{\text{ising}}$ as shown in Fig. 5c. a) Training and b) test loss for different qubit number $N$ as function of number of circuit parameters $M$, normalized by critical number of circuit parameters $M_{\text{c}}$. We have $L=2$ training states for $N=4$, and remarkably $L=1$ for $N=6,8$.
VIII.5 Stabilizer states
Finally, we apply the DQAE to stabilizer states. We generate training and test data as detailed in Sec. VIII.1, where we choose $V_{\text{Cliff}}$ to be a random Clifford circuit, and the input single-qubit states as one of the $6$ single-qubit stabilizer states, i.e. the eigenstates of the $\sigma^{x}$, $\sigma^{y}$ and $\sigma^{z}$ Pauli operators. The DQAE is trained with Algorithm 1. We show the test error for the DQAE after learning stabilizer states in Fig. 10. We find that the training with the algorithm nearly always converges, and the test error is only limited by the number of training states $L$. We numerically find that for at least up to $N\leq 10$ qubits the test errors collapse to a single curve when $C_{\text{test}}$ is normalized by the test error without training $C_{\text{test}}^{0}$ and we shift $L$ by $N/4$. This implies that the number of training states needed to generalize scales as
$L_{\text{c}}\propto N/4\,.$ (27)
Notably, the DQAE trained on stabilizer states is also able to disentangle non-stabilizer states of the form $V_{\text{Cliff}}\bigotimes_{j}|\phi_{j}\rangle$, where $|\phi_{j}\rangle$ are arbitrary (non-stabilizer) single-qubit states. This highlights that the DQAE is capable of out-of-distribution generalization [32, 28], i.e. it can generalize even when test and training data are drawn from different distributions.
Refer to captionRefer to caption
Figure 10: Test error $C_{\text{test}}$ for DQAE for stabilizer states as a function of number of training states $L$. DQAE is trained using Metropolis algorithm 1. We plot $C_{\text{test}}$ normalized to test error without training $C_{\text{test}}^{0}$.
IX Discussion
We introduced the DQAE to disentangle quantum states into single-qubit product states. Entangled states are highly susceptible to noise due to the collapse of the wavefunction, while product stares are more robust. This allows the DQAE to significantly enhance noise robustness of quantum information processing, such as the transfer of quantum information [51]. In particular, for the transfer of entangled states via a qubit loss channel, the DQAE gives an exponential advantage in the number of copies compared to unencoded transfer.
The DQAE can be trained with a simple cost function. We characterize overparameterization and generalization using the DQFIM, finding that for product states evolved with an Ising-type evolution, $M=O(N)$ circuit parameters and $L=1$ training data is sufficient for generalization. We also study DQAE for stabilizer states, finding that $L=O(N)$ training data are sufficient to generalize. Notably, we also showcase out-of-distribution generalization [32, 28], demonstrating that training on stabilizer states allows the DQAE to also disentangle a class of non-stabilizer states. Our algorithm for disentangling stabilizer states so far rests on a Metropolis type algorithm. We believe that a direct construction of the disentangling Clifford unitary can be found, similar to other dynamical learning problems for Cliffords [47, 52, 53].
Although we considered the DQAE for a qubit-loss model or (unrecoverable) leakage errors, it can also be applied to quantum error detection [2, 54]: In this case, $N$ logical qubits are redundantly encoded into many physical qubits. By performing check measurements, one can detect whether a logical error has occurred, and which qubit has been affected. Let us assume that each logical qubit has probability $p_{\text{L}}$ of experiencing an uncorrectable logical error. Without DQAE, one has to discard the full logical state if one logical qubit fails. In contrast, for states disentangled by DQAE, only the affected logical qubit has to be discarded, while the other logical qubits are unaffected. Thus, the DQAE could help to reduce the effective error rate for fault-tolerant quantum information transfer and storing information in quantum memory.
We note that the disentangling of quantum states could not only be used for noise suppression, but also for learning compact representations of quantum states [55]. Here, the DQAE can serve as a routine in quantum machine learning algorithms to learn simple representation of states.
Alternatively, the DQAE could be used to enhance cross-platform verification protocols [56]. In particular, measuring the fidelity $|\langle\psi_{1}|\psi_{2}\rangle|^{2}$ between two $N$-qubit quantum states $|\psi_{1}\rangle$, $|\psi_{2}\rangle$ on two different devices (without a quantum channel between them) requires exponential complexity $O(2^{N})$ in general [57]. Instead, one can first apply the DQAE $U$ on both states to disentangle the states into single-qubit product states $U|\psi_{1}\rangle$, $U|\psi_{2}\rangle$. Note that this leaves the fidelity invariant. Then, we can efficiently verify the fidelity between the states by local tomography of the $N$ single-qubit states which scales only as $O(N)$.
Finally, we comment on the relationship of DQAE to the recently proposed Clifford-enhanced matrix product state (CEMPS) [58] and their disentangling algorithms [58, 59, 60, 61]. Matrix product states (MPSs) are classical state representations with low entanglement, which can be combined with highly entangling Cliffords to be more expressible [58]. Clifford-disentangling of MPS is a classical simulation method that disentangles a (single) quantum state into a CEMPS. Such methods have applications for finding ground states [62] and simulating dynamics [63, 64, 65]. Notably, Clifford circuits doped with up to $N$ T-gates can be disentangled into a single-qubit product state with a Clifford unitary applied to it [59, 60]. We note that in contrast to Clifford-disentangling of MPS, DQAE is targeted to be trainable on quantum computers and can disentangle a large class of input states at the same time.
Future work could implement our DQAE in experiment. The DQAE can be straightforwardly implemented, following experimental demonstrations of other types of quantum autoencoders [18, 17, 15, 66].
Acknowledgements.
We thank Paul Bilokon, Michael Hanks and Joschka Roffe for enlightening discussions. This work is supported by the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (No. RS-2024-00413957) and the UK Hub in Quantum Computing and Simulation, part of the UK National Quantum Technologies Programme with funding from UKRI EPSRC grant EP/T001062/1.
Appendix A Increasing model capacity
In our discussion of the DQAE, we so far focused on mapping a dataset of $N$-qubit quantum states to $N$-qubit single-qubit product states. We can generalize this concept to involve a mapping from $N+K$ qubit states to an $N$-qubit product states, where $K\geq 0$. The additional $K$ qubits are mapped to a fixed $K$-qubit state $|a\rangle$, referred to as the reference state.
In the original quantum autoencoder architecture proposed by Romero et al. [7], the reference state $|a\rangle$ is typically set to $|0\rangle^{\otimes K}$. This state, having no further role in the protocol, is discarded after encoding the quantum state.
We propose an alternative approach, where instead of mapping the last $K$ qubits of every state in our dataset to a fixed reference state $|a\rangle$, we allow it to occupy a set of orthonormal states, such as the $K$-qubit computational basis states $\{|i\rangle\}_{i=1}^{2^{K}}$. As we will see, this enhances the capacity of the autoencoder. After measurement of the $K$ qubits, the orignal state can be reconstructed by a modified entangler: One prepares $K$ additional qubits which are initialized with the measurement outcomes $c_{k}$ of the disentangler. Refer to Fig. 11 for a sketch of the capacity-enhanced DQAE.
The training of the DQAE can be achieved by slightly modifying the original cost function of the DQAE: We choose $C_{\text{train}}^{\prime}=C_{\text{train}}+\sum_{k=1}^{K}\langle\psi|\sigma_{k+N}^{z}|\psi\rangle^{2}$, i.e. we add the square of the $\sigma^{z}_{k+N}$ measurements for each reference qubit. The cost function is minimized when the reference states are one of the computational basis state.
Figure 11: Illustration of an increased capacity DQAE that encodes a $N+K$-qubit state $|\psi\rangle$ into a product state comprising a $N$-qubit single-qubit product quantum state and $K$ classical bits $c_{k}=\{0,1\}$ by measuring the first $K$ qubits in the computational basis. The entangler $U^{\dagger}(\theta)$ takes as input computational basis states initialized with the $K$ measurement outcomes, along with the $N$ single-qubit product states, to reconstruct the original state $|\psi\rangle$.